A moving airplane dropped a cargo (ignore the effect of air resistance). If the plane from which the cargo was dropped was flying at a height of 600 m, and the cargo lands 300 m horizontally from the initial dropping point,


a. how fast was the plane flying when the cargo was released?

b. what was the velocity of the cargo right before it hit the ground?


1
Expert's answer
2022-03-10T11:15:38-0500

a. how fast was the plane flying when the cargo was released?

"l=v_0\sqrt{\frac{2h}{g}}"

"v_0=l/\sqrt{\frac{2h}{g}}\\
=300/\sqrt{\frac{2*600}{9.8}}=27.1\:\rm m/s"

b. what was the velocity of the cargo right before it hit the ground?

"v=\sqrt{v_0^2+2gh}\\
=\sqrt{27.1^2+2*9.8*600}=112\:\rm m/s"